### v1

@@
Priority-Neutral Matching Lattices Are Not Distributivethanks: We thank Ata Atay and Nicole Immorlica for helpful conversations.
Clayton Thomas Microsoft Research. — E-mail: clathomas@microsoft.com.
(April 2, 2024)
Stable matchings are a cornerstone of market design, with numerous practical deployments backed by a rich, theoretically-tractable structure. However, in school-choice problems, stable matchings are not Pareto optimal for the students. Priority-neutral matchings, introduced by [Ren22], generalizes the set of stable matchings by allowing for certain priority violations, and there is always a Pareto-optimal priority-neutral matching. Moreover, like stable matchings, the set of priority-neutral matchings forms a lattice [Ren22].
We study the structure of the priority-neutral lattice. Unfortunately, we show that much of the simplicity of the stable matching lattice does not hold for the priority-neutral lattice. In particular, we show that the priority-neutral lattice need not be distributive. Moreover, we show that the greatest lower bound of two matchings in the priority-neutral lattice need not be their student-by-student minimum, answering an open question of [Ren22]. This show that many widely-used properties of stable matchings fail for priority-neutral matchings; in particular, the set of priority-neutral matchings cannot be represented by via a partial ordering on a set of rotations. However, by proving a novel structural property of the set of priority-neutral matchings, we also show that not every lattice arises as a priority-neutral lattice, which suggests that the exact nature of the family of priority-neutral lattices may be subtle.
@@ 1 Introduction
Stable matchings enjoy a celebrated position within both the theory and practice of market design. A vast theoretical literature establish rich structural properties of stable matchings, and widespread practical deployments harness this theory to solve real-world problems.
@@
However, in some settings such as students-schools assignment, enforcing stability may be overly stringent, in two ways. First, stability may be overly stringent as a requirement. This is because students often have little ability to match outside of their centrally-assigned school; thus, priority violations—i.e., pairs $(d,h)$ where $d$ prefers $h$ to their match, and $h$ admits a student of lower priority than $d$ (or has an empty seat)—may be undesirable due to a sense of unfairness, but may not compromise the integrity of the matching market overall. Second, stability may be overly stringent in its implications. Namely, stability implies forgoing Pareto-optimality for the students. For instance, [APR09] report that in 2006 in New York City, over 4,000 grade 8 students could be made better off without worsening the match of any student. This leads to an important question: is there a natural alternative to stability which permits Pareto-optimal matches, while paying due respect to priority violations?111 The classical approach to Pareto-optimal matching in the school-choice problem is the Top Trading Cycles (TTC) mechanism [SS74, AS03]. While TTC is often a desirable mechanism, it (informally speaking) ignores priority violations completely, motivating the search for solution concepts that make a more integrated compromise between stability and Pareto-optimality.
Motivated by this question, [Ren22] proposes a generalization of stability, termed priority-neutrality. A matching $\mu$ is priority-neutral if, for each student $d$ whose priority is violated by $\mu$, it is not possible for another matching $\nu$ to make $d$ better off, unless $\nu$ also violates the priority of some student who is made worse off. Since stable matchings violate no student’s priority, each stable matching is priority-neutral. Additionally, priority-neutral matchings also allow for certain violations of $d$’s priority when, in some sense, $d$ cannot be improved without undue effects on others. [Ren22] proves that the set of priority-neutral matchings has mathematical structure: like the set of stable matchings, the set of priority-neutral matchings forms a lattice under the preferences of the students.222 See Section 2 for the definition of the ordering relation on matchings, and the definition of a lattice. Thus there is a unique student-optimal priority-neutral matching, which [Ren22] additionally shows is Pareto-efficient.333 [Ren22] is far from the first work to pursue the agenda of Pareto-improving stable matchings; see Section 1.1. Perhaps most notably, [EM20] introduce legal matchings, were a set $\mathsf{L}$ of matchings is legal if $\mathsf{L}=\{\mu\ \allowbreak|\ \allowbreak\text{$\forall\nu\in\mathsf{L}:$ \allowbreak no student $d$'s priority is violated by $\mu$ at school $\nu(d)$}\}$. [EM20] prove there is a unique legal set, and these legal matchings also form a lattice with a unique Pareto-efficient student-optimal outcome. However, the self-referential nature of the definition of a legal set motivates the search for other notions such as priority-neutrality, which [Ren22] posits may be a simpler definition than legality. Thus, known results suggest that priority-neutral matchings may be as theoretically tractable as stable matchings.
In this paper, we show that a crucial property of the set of stable matchings does not in fact carry over to priority-neutral matchings. In particular, we show that the priority-neutral lattice need not be distributive.444 A distributive lattice is one in which the equation $\mu\wedge(\nu\vee\xi)=(\mu\wedge\nu)\vee(\mu\wedge\xi)$ always holds, where $\wedge$ denotes greatest lower bounds and $\vee$ denotes least upper bounds. This lattice-theoretic property, while perhaps initially seemingly abstract, turns out to be intimately related to the theory of stable matchings. For instance, distributivity exactly characterizes stable matching lattices; i.e. a lattice arises as the set of stable matchings if and only if it is distributive [Knu74, Bla84, IL86]. Moreover, distributivity underpins the classical representation of the stable matching lattice by a partially ordered set of rotations [GI89] (since Birkhoff’s representation theorem [Bir37] says that distributive lattices are exactly those which can be represented as the downward-closed subsets of some partial order). This representation is widely exploited in both algorithmic and theoretical works on stable matchings (see, e.g., [Man13, EIV23]). Our result proves that the priority-neutral lattice cannot be represented by a partial order on some set of rotations.
Additionally, our construction resolves an open question of [Ren22]: we show that—unlike for stable matchings—the greatest lower bound of two matchings in the priority-neutral lattice need not be their student-by-student minimum. This result may be surprising: in contrast to stable matching lattices, as well as other well-known lattices from economic theory such as the lattice of Walrasian price vectors in an economy with gross substitutes consumer [GS99], lattice operations for priority-neutral matchings are not given by coordinate-by-coordinate minima / maxima.
This fact raises a natural question: what is the family of lattices that can arise as the set of priority-neutral matchings? We make progress towards this answer by proving that some lattices cannot be represented in this way, i.e., we show there exists a lattice $Q$ such that no set of preferences (and priorities) exist for which $Q$ is the priority-neutral lattice. Informally, our proof rules out certain lattices by showing that if there are many distinct priority-neutral matchings near the bottom (i.e., student-pessimal) priority-neutral matching, then there must also be many distinct priority-neutral matchings further above these elements. To complete the argument, we generalize the techniques of [Ren22] and exploit properties of other related matching lattices [EM20, FZ22] to prove a novel structural property of the priority-neutral lattice; for details, see Section 4.
In summary, we prove two main results. First, priority-neutral lattices fail to satisfy distributivity, a crucially important structural property of many lattices of economic importance. Second, not all lattices arise as priority-neutral lattices, showing that the precise nature of priority-neutral lattices may be subtle. Future works may give additional structural results clarifying how the priority-neutral lattice behaves. They may also look for alternative definitions which also reach efficient and fair matchings, according to appropriate new notions, with a more tractable structure.
1.1 Related Work
@@
This work primarily relates to two strands of literature: papers studying the structure of the set of stable matchings, and papers studying Pareto-improvements of stable matchings. Works on stable matchings follow from the seminal [GS62]. For structural properties specifically, [GI89] is the classical reference; see also [EIV23, Section 1.4]. Key individual works include [Knu74, Bla84, IL86], which establish that stable matching lattices are exactly characterized by distributive lattices. Other works establishing fundamental properties of stable matching mechanisms include [GS85a, MW71, Rot82, Rot84, DF81, GS85b].
@@
[Kes10] introduces the $\mathsf{EADAM}$ mechanism, which has become a centerpoint of the literature that seeks Pareto-improvements to stable matchings. [TY14] provides a simpler algorithm for computing $\mathsf{EADAM}$. [Ren22] introduces priority-neutral matchings, proved they form a lattice, and proved that the student-optimal matching in this lattice equals the outcome of $\mathsf{EADAM}$.
@@
Other than priority-neutral matchings, likely the most significant prior theoretical framework for Pareto-improving the set of stable matchings is the set of legal matchings [EM20], which we denote by $\mathsf{L}$; see footnote 3 for a definition. [Ren22] proves that every priority-neutral matching is legal (and that there can be legal, but not priority-neutral, matchings). [EM20] proves that $\mathsf{L}$ forms a lattice, and the student-optimal matching in $\mathsf{L}$ equals the outcome of $\mathsf{EADAM}$. Thus, these two matching lattices are closely related: the priority-neutral lattice is contained within the legal lattice, and they have the same student-optimal element. However, [EM20] show in fact that $\mathsf{L}$ forms a distributive lattice which is closed under taking the student-by-student minimum and maximum; in contrast, our main result demonstrates that the structure of priority-neutral lattice can be quite different from that of the legal lattice, since priority-neutral lattices need not satisfy these fundamental properties. Nevertheless, we exploit the connection between legal and priority-neutral matchings, along with additional properties of the legal lattice proven in [FZ22], in our structural result in Section 4.
@@
[EM20] follows the general idea of “stable sets” as in [NM44]; other papers taking similar or related approaches include [Ehl07, MVV11, AMV22, DE23], among others. [TDK20] define essentially stable matchings, which allow for Pareto-optimal matchings, but does not form a lattice (or characterize a unique Pareto-optimal matchings). [ACP+20, KS20, TZ21] compare mechanisms by the sets of priorities they violate. [CM23] studies a generalization of strategyproofness for $\mathsf{EADAM}$. [RTZ23] gives a novel characterization of $\mathsf{EADAM}$, as do many of the previously-mentioned papers.

### v2

@@
Characterization of Priority-Neutral Matching Latticesthanks: We thank Ata Atay, Nicole Immorlica, Phil Reny, participants at Stony Brook 2024, and various anonymous reviewers for helpful comments and conversations.
Clayton Thomas Yale University. — E-mail: thomas.clay95@gmail.com.
(December 7, 2025)
We study the structure of the set of priority-neutral matchings. These matchings, introduced by [Ren22], generalize stable matchings by allowing for priority violations in a principled way that enables Pareto-improvements to stable matchings. Known results show that the set of priority-neutral matchings is a lattice, suggesting that these matchings may enjoy the same tractable theoretical structure as stable matchings.
In this paper, we characterize priority-neutral matching lattices, and show that their structure is considerably more intricate than that of stable matching lattices. To begin, we show priority-neutral lattices are not distributive, an important property that characterizes stable lattices and is satisfied by many other lattice structures considered in matching theory and algorithm design. Then, in our main result, we show that priority-neutral lattices are in fact characterized by a more-involved property which we term being a “movement lattice,” which allows for significant departures from the order theoretic properties of distributive (and hence stable) lattices. While our results show that priority-neutrality is more intricate than stability, they also establish tractable properties. Indeed, as a corollary of our main result, we obtain the first known polynomial-time algorithm for checking whether a given matching is priority-neutral.
@@ 1 Introduction
Stable matchings are a foundational notion in the theory of matching under preferences. Since their introduction in [GS62], they have been found to exhibit a both textbook algorithmic elegance and rich theoretical structure [Knu76, Bla84, GI89, KT06]. These insights culminate in a striking characterization: when matchings are ordered according to agents’ preferences, the set of stable matchings forms a distributive lattice. [^1] Specifically, matchings are partially ordered according to preferences as follows: if we are matching students to schools, $\mu\geq\nu$ means that for all students $d$, we have $\mu(d)\succeq_{d}\nu(d)$. A lattice is a partially-ordered set such that each pair $x,y$ has a greatest lower bound $x\wedge y$ and a least upper bound $x\vee y$. A distributive lattice is one in which the equation $\mu\wedge(\nu\vee\xi)=(\mu\wedge\nu)\vee(\mu\wedge\xi)$ always holds. Moreover, every distributive lattice arises in this way, so this property precisely characterizes the structure of stable matchings. This characterization has been instrumental for countless results on stable matchings, including their combinatorics (e.g., [KGW18, PP22]), communication (e.g., [GNOR19, ABKS20, GT24]), and algorithms (e.g., [Man13, EIV23]).
@@
Despite these successes, there are compelling reasons to look beyond stable matchings. For instance, in the crucial real-world setting of student-to-school assignment [AS03, APRS05, APR05, Sön23], stability may not be a binding requirement, and is often overly restrictive. First, regardless of the schools’ priorities, students are typically unable to attend a school that the centralized algorithm did not assign them to. Thus, priority violations in matching $\mu$—i.e., pairs $(d,h)$ where $h\succ_{d}\mu(d)$ and $d\succ_{h}\mu(h)$—may create a sense of unfairness, but may not compromise the integrity of the matching overall. Second, stability implies forgoing Pareto-optimality for the students. For instance, [APR09] report that in 2006 in New York City, over 4,000 grade 8 students could be made better off without worsening the match of any student. This leads to an important question: is there a natural class of matchings which permits Pareto-improvements to stable matchings, while paying due respect to priority violations? [^2] The classical Pareto-optimal matching mechanism in student-to-school assignment is Top Trading Cycles (TTC) [SS74, AS03]. Despite having many desirable properties, TTC (informally speaking) ignores priority violations completely, motivating the search for a more integrated compromise between stability and Pareto-optimality.
Motivated by this question, the recent breakthrough work [Ren22] proposes a generalization of stability, termed priority-neutrality. When matching applicants to institutions, a matching $\mu$ is priority-neutral if, for each applicant $d$ whose priority is violated by $\mu$, it is not possible for another matching $\nu$ to make $d$ better off, unless $\nu$ also violates the priority of some applicant who is made worse off. Since stable matchings violate no applicant’s priority, every stable matching is priority-neutral. But unlike stable matchings, priority-neutral matchings also allow for violations of $d$’s priority when, in some sense, $d$ cannot be improved without undue effects on others.
[…]
Our Results.
In this paper, we show that the structure of the priority-neutral lattice is in fact far more involved than that of the stable lattice. Our first result (Section 3) begins by answering an open question of [Ren22], and showing that the priority-neutral lattice is not distributive. Moreover, we show that in the priority-neutral lattice, the greatest lower bound of two matchings is not their coordinatewise minimum (as it is in the stable matching lattice, and in many other lattices of algorithmic interest, e.g., [PQ80, KNK93, GS99]). [^4] The coordinatewise minimum of matchings $\mu$ and $\nu$ pairs each applicant $d$ to their worse choice from $\{\mu(d),\nu(d)\}$. In the stable matching lattice, $\mu\wedge\nu$ is the coordinatewise minimum of $\mu$ and $\nu$, and $\mu\vee\nu$ is (the analogous notion of) the coordinatewise maximum of $\mu$ and $\nu$. One can show that, if $\wedge$ and $\vee$ are given by coordinatewise minima and maxima (respectively), then the lattice is distributive. We prove that in the priority-neutral lattice, $\wedge$ is not given by coordinatewise minima (although, as shown by [Ren22], $\vee$ is given by coordinatewise maxima). These facts imply that priority-neutral matchings lack much of the tractable structure of stable matchings. (For instance, they show that the priority-neutral lattice cannot be represented by a partial order on a set of rotations, the classic strengthening of Birkhoff’s representation theorem [Bir37] for the stable lattice which underpins the results discussed above.)
Given that priority-neutral lattices thus lack the tractable structure of stable matching lattices, we ask the following questions: What lattices do arise as priority-neutral matching lattices? What properties to priority-neutral lattices satisfy, and how can we represent their structure?
In our main result (Section 5), we characterize priority-neutral matching lattices. That is, we provide a representation of the set of priority-neutral matchings, and moreover, we delineate the exact class of lattices which can arise in this way. This class—which we call movement lattices—is fairly intricate, and we show via examples that priority-neutral lattices can have order-theoretic structure quite different from that of distributive lattices. Moreover, proving that this representation captures priority-neutral matching lattices is involved: we build on an extensive theory of different matching lattices that generalize stability, and use techniques from [Kes10, TY14, EM20, Ren22, FZ22]. For some preliminary intuition, see below.
While our results imply that priority-neutral lattices are more intricate than stable ones, they also show ways in which priority-neutrality is tractable. For instance, our characterization directly yields a poly-time algorithm for checking whether a given matching is priority-neutral. Prior to our work, it was far from clear how to efficiently check whether a matching $\mu$ is priority-neutral, since the definition of priority-neutrality requires verifying that no alternative matching $\nu$ satisfies some condition (specifically, $\nu$ cannot improve the match of an applicant whose priority is violated in $\mu$ without violating the priority of an applicant who is worse off in $\nu$).
Intuition for Our Results.
Intuitively, the structure of the stable matching lattice is derived from the fact that avoiding priority violations—i.e., pairs $(d,h)$ with $h\succ_{d}\mu(d)$ and $d\succ_{h}\mu(h)$—enforces simple ordering-based constraints on the set of stable matchings. For instance, in any stable matching $\mu$ which sets $\mu(h)=d_{0}$, consider any potential priority-violating pair $(d,h)$ with $d\succ_{h}d_{0}$. Then, to preserve stability, we must guarantee that $\mu$ matches $d$ at least as well as $h$, i.e., $\mu(d)\succ_{d}h$. Thus, this potential priority violation gives a natural ordering structure on the differences between stable matchings: any time $h$ is made worse-off than $d$, we must in turn make $d$ better-off than $h$.
For a simple example of the above discussion, consider Figure 1. In this instance, we cannot assign $h_{D}$ to $d_{C}$ without “first” assigning $d_{A}$ to $h_{B}$ (or else $d_{A}$’s priority would be violated at $h_{D}$). More generally, all stable matching lattices can be represented this way, where different changes can be made to the matching so long as all of their “predecessor” changes have already been made. (The idea of such predecessor relations exactly characterizes distributive lattices by Birkhoff’s representation theorem [Bir37], and the formal notion capturing this for stable matchings is the rotation DAG; see Section 4.1.1.555In Appendix A, we show for the sake of exposition that Figure 1 can be directly generalized to construct a stable matching lattice isomorphic to any distributive lattice. We believe our construction simplifies prior ones showing this fact [Bla84, GI89]. )
Institutions’ Priorities:
$h_{A}$ $h_{B}$ $h_{C}$ $h_{D}$
$d_{A}$ $d_{B}$ $d_{C}$ $d_{D}$
$d_{B}$ $d_{A}$ $d_{D}$ $d_{A}$
$d_{C}$
Applicants’ Preferences:
$d_{A}$ $d_{B}$ $d_{C}$ $d_{D}$
$h_{B}$ $h_{A}$ $h_{D}$ $h_{C}$
$h_{D}$ $h_{B}$ $h_{C}$ $h_{D}$
$h_{A}$
$\mu_{1}$ $=\{(h_{A},d_{A}),(h_{B},d_{B}),(h_{C},d_{C}),(h_{D},d_{D})\}$
$\mu_{2}$ $=\{(h_{A},d_{B}),(h_{B},d_{A}),(h_{C},d_{C}),(h_{D},d_{D})\}$
$\mu_{3}$ $=\{(h_{A},d_{B}),(h_{B},d_{A}),(h_{C},d_{D}),(h_{D},d_{C})\}$
Figure 1: An example instance where the stable matching lattice is $\{\mu_{1},\mu_{2},\mu_{3}\}$, with $\mu_{1}<\mu_{2}<\mu_{3}$. This illustrates the simple ordering constraints on stable matchings: we must ensure that $d_{A}$ is improved “before” $h_{D}$ is made worse. In contrast, we show that the ordering constraints on priority-neutral matchings are more involved: they may require only that some change, among many possibilities, is made “before” another change.
Where stable matchings are those that do not have priority violations, priority-neutral matchings are those matchings $\mu$ that do not have what we call priority-correcting adjustments—i.e., alternative matchings $\nu$ that have $\nu(d)\succ_{d}\mu(d)$ for some $d$ whose priority is violated in $\mu$, and such that there does not exist a $d_{x}$ whose priority is violated in $\nu$ and $\mu(d)\succ_{d_{x}}\nu(d)$. According to [Ren22], such matchings $\nu$ provide a plausible argument for such a applicant $d$ as to why their priority should not have been violated in matching $\mu$, namely, $\nu$ “corrects” $\mu$ from $d$’s perspective by improving applicant $d$’s match, while only harming applicants $d_{x}$ if their priorities are respected.
Unfortunately, avoiding priority-correcting adjustments does not enforce the same simple ordering structure on the set of priority-neutral matchings (compared to the simple structure that avoiding priority violations enforces on the set of stable matchings). Intuitively, this is because if $\mu$ violates $d$’s priority, and if a potential priority-correcting adjustment $\nu$ violates the priority of multiple applicants $d_{x}$ and $d_{y}$, then this $\nu$ can witness the fact that $\mu$ is not priority-neutral only if neither $d_{x}$ or $d_{y}$ have been made better off than their match in $\nu$. Hence, the change that violates $d$’s priority can be made after either $d_{x}$ or $d_{y}$ have been improved. The resulting ordering structure can be much more intricate than those that (by Birkhoff’s theorem) underpin stable and distributive lattices (which could only require, for example, that both $d_{x}$ and $d_{y}$ must have been improved).
In our first result (in Section 3), we construct an instance implementing the above idea for priority-neutral lattices, showing that they are not distributive. After this (in Section 4), we directly pursue a generalization of the above discussion, and attempt to capture the order-theoretic constraints induced by priority-correcting adjustments in terms of a “multiple predecessor” generalization of results for stable lattices. While this approach allows us to prove some structural results on the set of priority neutral matchings, it does not give a tight characterization (see Section 4.3.4 for discussion). Hence, in our main result (in Section 5), we take a more “low level” approach, and encode the data of a priority-correcting adjustment into a direct representation. In the end, this low-level approach gives a tight characterization.
Related Work.
@@
Our paper builds most directly on two strands of literature. First, we build on a vast literature on the structure of stable matchings following from the seminal [GS62]. For structural properties specifically, [GI89] is the classical reference; see also [EIV23, Section 1.4] or [CT19]. Key individual works include [Knu74, Bla84, IL86], which establish that stable matching lattices are exactly characterized by distributive lattices. Other works establishing fundamental properties of stable matching mechanisms include, for example, [GS85a, MW71, Rot82, Rot84, DF81, GS85b].
@@
Second, we build on papers studying Pareto-improvements of stable matchings. [Kes10] introduces a mechanism known as $\mathsf{EADAM}$, which has become a centerpoint of this literature. [TY14] provides a simpler algorithm for computing $\mathsf{EADAM}$. [Ren22] introduces priority-neutral matchings, proves they form a lattice, and proves that the applicant-optimal matching in this lattice equals the outcome of $\mathsf{EADAM}$.
@@
Other than priority-neutral matchings, likely the most significant prior theoretical framework for Pareto-improving the set of stable matchings is the set of legal matchings [EM20], which we denote by $\mathsf{L}$; see Footnote 3 for a definition. [EM20] proves that $\mathsf{L}$ forms a lattice, and the applicant-optimal matching in $\mathsf{L}$ equals the outcome of $\mathsf{EADAM}$. [Ren22] proves that every priority-neutral matching is legal (and that there can be legal, but not priority-neutral, matchings). Thus, these two matching lattices are closely related. Throughout this paper, we exploit the connection between priority-neutrality and legality, along with additional structural results on legality proven by [FZ22] (for more background, see Section 4.1). However, we note that [EM20] show that $\mathsf{L}$ forms a distributive lattice which is closed under coordinatewise minimum and maximum operations. In contrast, we prove the priority-neutral lattice is not distributive, and hence is quite different in structure from the legal or stable lattice.
@@
Many variations on stable matching have been proposed. [EM20] follows the general idea of “stable sets” as in [NM44]; other papers taking similar or related approaches include [Ehl07, MVV11, AMV22, DE23], among others. [TDK20] define essentially stable matchings, which allow for Pareto-optimal matchings, but does not form a lattice (or characterize a unique Pareto-optimal matchings). [ACP+20, KS20, TZ21] compare mechanisms by the sets of priorities they violate. [CM23] studies a generalization of strategyproofness satisfied by $\mathsf{EADAM}$. [RTZ23] gives a novel characterization of $\mathsf{EADAM}$, as do many of the previously-mentioned papers.
In certain more-general matching problems, stable matchings are known to form a non-distributive lattice. Specifically, this is the case in many-to-many [Bla88] and many-to-one [EO04] matching settings with substitutes preferences; in fact, the recent paper [EF25] shows that all finite lattices can arise as the stable lattice in this model. These facts crucially use the generality of multi-partner matching with substitutes in preferences; indeed, when preferences are responsive (i.e., institutions preferences are consistent with a preference list over applicants) [Rot86], or when they satisfy the commonly-assumed law of aggregate demand [Fle03, HM05], stable matchings form a distributive lattice [Alk02, Li14]. We believe it is very interesting that priority-neutral lattices have a non-distributive and fundamentally new structure, despite being defined in the completely standard one-to-one matching model involving no new primitives.
Following [Knu74, GI89], lattice properties of stable matchings have been instrumental in many results. [KGW18, PP22] make use the rotation poset to make breakthrough progress on the combinatorial problem of determining the maximum possible number of distinct stable matchings as a function of $n$, the number of applicants and institutions. [GT24] designs a blackboard verification (in the sense of NP) communication protocol for the applicant-optimal stable mechanism. Much algorithmic work on different variants of such matching problems also exploits rotation poset representations and distributive lattice properties, see e.g. [Man13, EIV23]. Additionally, [IL86, SS15] use related arguments to show that counting stable matchings, and other combinatorial matching problems, are $\mathsf{P}-$hard, and [Sub94, CFL14] show that stable matchings are connected to a circuit model with “comparator” gates, and to certain novel complexity classes between $\mathsf{NL}$ and $\mathsf{P}$. Further from our work, [Pit89, IM05, AKL17, GMM21, KMQ21, Mau21, CT22] study stable matchings under certain random preference distributions. Additionally, different (non-stable) matchings have been defined using ordinal preferences. One example is popular matchings following [AIKM07]; the structural properties of stable matchings have found applications for this set of matchings as well [Kav24]. Another line of work investigates the extent to which natural decentralized matching dynamics converge to stable matchings [RV90, AGM+08, Rud24]. [^6] To our knowledge, no such decentralized dynamics have been investigate that accommodate Pareto-optimal matchings or Pareto-improvements to stable matchings; this may be an interesting topic for future work.
Paper Organization.
We recall the definition of stability and priority-neutrality in Section 2, and show that the priority-neutral lattice is not distributive in Section 3. In Section 4, we build up to our main characterization as follows: first, we recall additional preliminaries and results from prior work; second, we establish new structural properties of the priority-neutral lattice; and third, we provide a simpler sufficient representation of priority-neutral lattices (i.e., a representation which includes all priority-neutral lattices, but which includes other lattices as well). In Section 5, give our main characterization result: a necessary and sufficient representation of priority-neutral lattices. As a corollary, we also provide the first poly-time algorithm for checking whether a matching is priority-neutral.